OpenStudy (anonymous):

some help please, 7+3(2a-5)-6a-8b)-2b

OpenStudy (anonymous):

put val;ues in this: (2b + 5) - (8b + 6) + 3 (b - 2) A - is not solely a -. It's a -1. You multiply the number outside the bracket with every number inside the bracket. = 2b + 5 - 1 (8b + 6) + 3 (b - 2) = 2b + 5 - 1 * 8b - 1 * 6 + 3 * b - 3 * 2 = 2b + 5 - 8b - 6 + 3b - 6 = 2b - 8b + 3b + 5 - 6 - 6 = -3b - 7. I hope that helps. :)

OpenStudy (lgbasallote):

according to PEMDAS...first thing to do is multiplication so let us multiply via distributive property... 7 + 3(2a) + 3(-5) - 6a - 8b -2b would you clarify where that parenthesis is after 8b? what does it group exactly??

OpenStudy (anonymous):

thank you Rohangrr. yes helped very much